OpenStudy (anonymous):

The steps in the simplification of the expression 6 + 3x + 4y − 2x – y − 7 are shown below. Step 1: 6 – 7 + 3x − 2x + 4y − y Step 2: (6 − 7) + (3x − 2x) + (4y − y) Step 3: −1 + x (3 − 2) + y (4 − 1) Step 4: −1 + x (1) + y (3) Step 5: −1 + x + 3y Which statement is correct? Step 2 uses associative property of addition, and step 3 uses distributive property of addition. Step 1 uses commutative property of addition, and step 5 uses associative property of addition. Step 4 uses associative property of addition, and step 3 uses commutative property of additi

mathslover (mathslover):

I think the second step is right just note there that all the terms are re arranged : that is the main thing about commutative property : of addition a+b=b+a

OpenStudy (anonymous):

Oh okay. I see what you mean.

mathslover (mathslover):

oops sorry i am wrong wait

OpenStudy (anonymous):

lol how?

mathslover (mathslover):

2nd option states that step 5 uses associative property that is not true

mathslover (mathslover):

1st option is right

OpenStudy (anonymous):

ok. and how is it distributive on the third statmenet

mathslover (mathslover):

we took common : ab+ac=a(b+c)